OpenStudy (anonymous):

30 mi / h = ____ ft / min Please help.

OpenStudy (anonymous):

a. 2640 b. 264 c. 880 d. 2112

OpenStudy (anonymous):

@Hero please hurry and reply...

hero (hero):

\[\frac{30 \text{mi}}{\text{h}} = \frac{x\space\text{ft}}{\text{min}}\] \[\frac{30 \text{(5280) ft}}{\text{60 min}} = \frac{x\space\text{ft}}{\text{min}}\] \[\frac{30 \text{(5280) ft}}{\text{60 min}}\dot\ \frac{\text{min}}{\text{ft}} = x\] \[\frac{30 \text{(5280)ft}}{\text{60 min}}\dot\ \frac{\text{min}}{\text{ft}} = x\]

hero (hero):

\[\frac{30(5280)}{60} = x\]

OpenStudy (anonymous):

So the answer would be C? 880?

OpenStudy (anonymous):

NO!

OpenStudy (anonymous):

ITS A

hero (hero):

No. How did you get that?

hero (hero):

I know it is.

hero (hero):

What did you do wrong?

hero (hero):

No one told you to pick c

hero (hero):

\[\frac{30(5280)}{60} = 2640\]

OpenStudy (anonymous):

Because 30 x 5280 divided by 60 would equal 2640. Instead of multiplying I just divided 5280 into 60. Like a dummy. & Excuse me.. I made a mistake...

OpenStudy (anonymous):

Okay thank you :) bye.

OpenStudy (anonymous):

Ahaaha, you had to use edgenuity too, i see. XD